A canister of butane gas (C4H10, MM=58.12 g/mol) used for portable cooking is available in a store in two sizes, 750 mL and 2.0 L.  The 750 mL canister contains 85 g of butane gas with a pressure of 48 atm at 28 C.  How many grams of butane gas are present in the bigger canister under the same condition of pressure and temperature?


1
Expert's answer
2022-04-26T10:01:02-0400

V1/V2 = m1/m2

m= m1V/ V1

m2 = (85 g * 2 L) / 0.75 L = 226 g


or


PV = nRT

n2 = ( 48 * 2) / (0.8206 * 301.15) = 3.88 moles

m2 = 3.88 * 58.12 = 226 g